2018 Operation

The place to discuss the LHC. Commissioning, operation, issues, events ....
Post Reply
dukwon
Posts: 44
Joined: Fri Apr 10, 2015 1:04 pm

2018 Operation

Post by dukwon » Sun Mar 11, 2018 12:33 am

The LHC has been handed over to OP, and powering tests should begin Monday.

If you haven't done so already, you can catch up with the slides from Chamonix while waiting for first beams: https://indico.cern.ch/event/676124/tim ... l.detailed

Here's a nice article from the Bulletin:
https://home.cern/cern-people/updates/2 ... ibernation

To summarise:
  • 60 /fb for ATLAS and CMS
  • 2 /fb for LHCb
  • 131 days of physics
  • 4 weeks of heavy ions
  • 1 week of desqueezed beams
  • Levelling at IP1&5 to a pile-up of 55 to 60
  • 25—27 cm β* with ATS
  • 20 days of MD
Here's hoping for a nice physics production year to round off Run 2.

dukwon
Posts: 44
Joined: Fri Apr 10, 2015 1:04 pm

Re: 2018 Operation

Post by dukwon » Tue Mar 20, 2018 10:04 am

The short-term schedule is currently being driven by the intervention on the CMS pixel detector.

Of course things are subject to change, but here are some planned milestones

22nd March: first dry run with handshakes and mode changes
27th March: TED shots
29th March: vacuum valves open around CMS, possibly test of beam and injection permits
1st April: first beams
23rd April: first stable beams

jmc2000
Posts: 147
Joined: Sun May 08, 2011 2:51 pm

Re: 2018 Operation

Post by jmc2000 » Thu Mar 29, 2018 9:13 pm

Late February LHCC meeting:
https://indico.cern.ch/event/692482/con ... Status.pdf

Looks as if the number of days for the p-p run has been reduced from 131 to 124 to accommodate special runs:
https://beams.web.cern.ch/sites/beams.w ... e_2018.pdf

jmc2000

dukwon
Posts: 44
Joined: Fri Apr 10, 2015 1:04 pm

Re: 2018 Operation

Post by dukwon » Fri Mar 30, 2018 10:45 am

Both beams successfully circulating

https://i.imgur.com/3ef5c2g.png

dukwon
Posts: 44
Joined: Fri Apr 10, 2015 1:04 pm

Re: 2018 Operation

Post by dukwon » Tue Apr 10, 2018 10:02 am

The recommissioning is going very smoothly. Collision-finding should already take place this week, and first stable beams with 3 bunches could be as early as Monday 16th.

There will be a scrubbing run either next week or the one after.

dukwon
Posts: 44
Joined: Fri Apr 10, 2015 1:04 pm

Re: 2018 Operation

Post by dukwon » Wed Apr 11, 2018 7:04 pm

First collisions (not stable beams) could be tomorrow mid-morning if loss maps and TCDI validation are performed during the night.

User avatar
DCWhitworth
LHCPortal Guru
Posts: 568
Joined: Mon Nov 30, 2009 8:13 am
Location: Norwich, UK

Re: 2018 Operation

Post by DCWhitworth » Thu Apr 12, 2018 7:28 am

Thanks for keeping us all updated. I think I shall be having Page 1 and the LHC Announcer open on my work PC today :-)
DC

The LHC - One ring to rule them all !

dukwon
Posts: 44
Joined: Fri Apr 10, 2015 1:04 pm

Re: 2018 Operation

Post by dukwon » Thu Apr 12, 2018 9:22 am

dashboard0.png
We have collisions!

dukwon
Posts: 44
Joined: Fri Apr 10, 2015 1:04 pm

Re: 2018 Operation

Post by dukwon » Sat Apr 14, 2018 11:21 am

ADT tests (necessary for train injections and abort gap cleaning) have been given priority over first stable beams w/ 3b, so the latter has been pushed back to Tuesday.

dukwon
Posts: 44
Joined: Fri Apr 10, 2015 1:04 pm

Re: 2018 Operation

Post by dukwon » Mon Apr 16, 2018 6:47 pm

Latest news is that we could see stable beams after tomorrow's morning meeting (so in about 12 hours or more)

dukwon
Posts: 44
Joined: Fri Apr 10, 2015 1:04 pm

Re: 2018 Operation

Post by dukwon » Tue Apr 17, 2018 11:02 am

Stable beams with 3 bunches per beam!
lhc1.png

dukwon
Posts: 44
Joined: Fri Apr 10, 2015 1:04 pm

Re: 2018 Operation

Post by dukwon » Tue Apr 17, 2018 11:10 am

Screenshot from 2018-04-17 13-07-37.png
3DTower.png
Screenshot from 2018-04-17 13-07-00.png

dukwon
Posts: 44
Joined: Fri Apr 10, 2015 1:04 pm

Re: 2018 Operation

Post by dukwon » Sat Apr 21, 2018 11:37 am

We've had 3 fills with 75b, where the third was with BCMS beams. I believe the next step will be 300b.
Currently scrubbing with 2605b, 2 days earlier than planned. Even the slides from this morning said scrubbing on Monday.

Post Reply